OpenStudy (anonymous):

An electron having an initial velocity of 10*3 cm/s is directed from a distance of 1mm at another electron whose position is fixed.How close to the stationery electron will the othere approach.

OpenStudy (jamesj):

Work = Force.Distance. You to find the work done on the moving electron until the Kinetic energy is zero. In other words, find the distance x such that \[ \Delta KE = \frac{1}{2}m_ev_0^2 - 0 = \int_{1 \ mm}^{x} \frac{ke^2}{r^2}dr \]